# Preserving Multilingual Quality While Tuning Query Encoder on English Only

### v1

@@ Appendix A Triplets from ARXIV subset
Table 12: Evaluations of the $L12$ query encoder tuned with a frozen embedding block, learning rate 5e-8, batch size 14 and different triplet loss margins (first column). Values that did not pass the two-tailed test are shown in gray.
@@ B Sch D cdcdcdc+/- d+/- c+/- d+/-
Appendix B Dependency on tuning parameters
B.1 Learning rate
Our results (Section 3, Table 2) do not depend strongly on the tuning parameters, but an order of magnitude change in learning rate does have some effect. As shown in Table 3, the rates in the range 2e-8 to 6e-8 provide similar results; the lower rates 2e-8 to 4e-8 may be slightly safer on languages, with weaker gain on English data. The rate 1e-8 is too low: tuning barely made any changes. The higher rates 7e-8 to 1e-7 give higher gains on English data, but the number of degraded language pairs for cosine matching jumps from 1-2 to around 20-25 (out of 225).
B.2 Batch size
In Table 4 we show results for batch sizes 7, 14, 28, 56 and 112, while keeping the number of samples per epoch the same (14000). The row with batch 14 here coincides with row with learning rate 5e-8 in Table 3, and with row for frozen embedding block in Table 2. The results for all batch sizes are similar. Tuning with the higher batch size 112 is a bit ’safer’ for languages, not degrading any language pair when evaluated by cosine measure, and degrading only one language pair (for entailment vs contradiction) when evaluated by distance measure. This comes at the price of lower gains on MSMARCO and ARXIV.
Table 5 shows what happens if the number of batches per epoch (1000) is kept the same, rather than the number of samples. In this setting the larger batch size 112 leads at tuning to a less frequent validation (by MSMARCO validation subset) and, effectively, to later and less reasonable stopping. This results in higher gain on MSMARCO test subset, but in far worse results on ARXIV and XNLI.
B.3 Margin of triple loss
The margin is an important parameter of the contrastive learning through the triplet loss. In Table 6 we show the dependency of the evaluation results on the margin during its tuning. We consider our default tuning parameters (Section 2.3), but change the margin. The results are natural: a margin up to 0.15 is reasonable, and at higher margins the disturbance on cross-lingual, and, eventually, on English data evaluation gets too strong.
B.4 Stopping criterion
In Table 7 we show dependency of improvement results on the stopping criterion. The stoppings after 5 or 10 non-improvement epochs give similar results. Stopping after 15 non-improvement epochs continues the trend of increased gain on English data, but with a deterioration for a few language pairs.
Appendix C L12 Model
For comparison with our consideration (Section 3) of the $E5$ model, here we show a few results of the query model tuning (in the same regime, see Section 2.3) that instead starts with the $L12$ model. Altogether, we observe that $L12$ does not preserve languages as nicely as $E5$, even for the cosine measure. And yet the tuning of $L12$ makes much more language pairs improved than degraded. While each language pair is important in its own right, we find it interesting that even for $L12$ the change in the absolute number of errors summed up across all language pairs is still within $1\%$.
C.1 Freezing layers
C.2 Learning rate
In Table 9 we show evaluations of $L12$ model tuned at different learning rates. This table is arranged similarly to the $E5$ model Table 3. The results for $L12$ are not as good as for $E5$, this is especially noticeable by the number of language pairs that have degraded.
C.3 Batch size
Dependency of $L12$ tuning on the batch size is shown in Table 10 (number of samples per epoch is 14000) and in Table 11 (number of batches per epoch is 1000). Observations are somewhat similar to $E5$ (Appendix B.2, except that generally $L12$ performs not as good as $E5$ and the batch size 7 turns out bad for $L12$.
C.4 Margin of triple loss
Dependency of $L12$ tuning on the contrastive learning margin size is shown in Table 12. The results for margin 0.1 are distinctly better. Altogether, the model $L12$ appears to be more sensitive (compared to $E5$) to the tuning parameters if the goal is to preserve performance on multilingual data XNLI and on out-of-domain data ARXIV.
Appendix D Effects of Learning Rate Scheduler and Weight Decay
Using the fine-tuned E5 model with the frozen embedding block, tuned using a batch size of 14, learning rate of $10^{-7}$ and a margin of 0.1, we randomly vary the batch size, learning rate scheduler and weight decay in order to assess their impact on the model’s final performance. In Table 13, we present the change in performance across these different configurations. Table 14 lists the different schedulers we considered. Values in blue indicate the top improvements whereas values in red indicate the worse degredations.
Across these parameters, we see that, on average, the batch size appears to have to most significant impact, generally leading to poorer performance as the batch size is decreased. Within each batch size group, we see that using an exponential learning rate scheduler($E_{0.95}$ or $E_{0.98}$) is generally worse than using any of the other schedulers or no scheduler at all. A specific exception exists when using a batch size of 100 where the exponential scheduler outperforms the quadratic one.
@@ Appendix D Effects of Learning Rate Scheduler and Weight Decay
Appendix E Improvement on XNLI, by language pairs
In Table 2 and in other similar tables we have been showing the improvement on XNLI dataset by a number of improved or degraded language pairs, when the changes are statistically significant.
Figure 1: Improvement of $E5$ on XNLI entailment-neutral comparisons assessed by cosine, shown by language pairs. The improvement results from tuning on MSMARCO with embedding block frozen and tuning parameters as specified in Section 2.3.
Figure 2: Improvement of $E5$ on XNLI entailment-contradiction comparisons assessed by cosine.
Figure 3: Improvement of $L12$ on XNLI entailment-neutral comparisons assessed by cosine.
Figure 4: Improvement of $L12$ on XNLI entailment-contradiction comparisons assessed by cosine.
Here we have shown in Figures 123 and  4 the actual improvement calculated for each language pair separately, by equation 1, using the count of errors for the language pair on XNLI.

### v3

@@ Appendix I Significance Test
Table 12: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, learning rate 1e-7, batch size 56, margin 0.1 and a range of weight decay (first column). Values that did not pass the two-tailed test are shown in gray.
weight msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
decay cdcdcdc+/- d+/- c+/- d+/-
5 6.53 7.26 3.81 4.50 3.51 5.96 222/0 216/0 197/2 150/14
1 7.26 8.73 4.77 5.39 3.76 7.95 222/0 219/0 201/2 160/20
0.5 7.30 8.82 4.90 5.39 3.51 7.73 222/0 217/1 201/2 158/20
0.1 7.30 8.82 4.90 5.39 3.51 7.73 222/0 217/1 201/2 158/20
Table 13: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, learning rate 5e-8, batch size 14, margin 0.1 and a range of weight decay (first column). Values that did not pass the two-tailed test are shown in gray.
A weight decay may restrict increase of model weights, but it does not improve the evaluation results. We show some representative results in Tables 12 and 13. While restricting gains on the tuning goal, weight decay does not help to preserve the other qualities: the results on XNLI and ARXIV are no better than without weight decay. If there is any recipe for further improving the gains both on the tuning goal and on the related qualities, it has to be a less crude interference into the tuning.
Since weight decay may be more effective at higher learning rates, the parameters for Table 12 are chosen at higher rate and batch size, compared to our ’default’ choice, which is used in Table 13. The learning rates and batch sizes of these tables relate by square root scaling rule (see Section M.1.1).
M.3 Candidate layers for freezing
Figure 11: Evaluations of the $E5$ query encoder tuned with a frozen embedding block and all layers ‘out-put.dense.weight’, with batch size 14, margin 0.1 using different learning rates on (a) XNLI and (b) the English-only datasets (MSMARCO and ARXIV). Values that did not pass the two-tailed test are shown with open markers.
From evaluation results in Figure 2 we may consider the learning rate below 7e-8 (but above 1e-8) as safely suitable for adiabatic tuning. But we know this only because we evaluated the tuned models on the out-of-tuning domains ARXIV and XNLI.
Is there any way to know the upper boundary without having extensive data for evaluation? Could there be an empirical recommendation not to exceed certain learning rate? Can we increase the adiabatic tuning range of learning rate?
In attempting to answer these questions, we considered the largest changes in the layers at different learning rates. One suspect layer, by simple crude measures, is output.dense.weight. Here in Tables 14 and 15 we show why we selected this layer. Our motivation here is based on a simple and crude criteria; more detailed research and understanding may reveal better ways to extend the adiabatic tuning regime.
The gains from the tuning by freezing the layer output.dense.weight (in each transformer block) are shown in Figure 11. In comparison to the default tuning (Figure 2) we can see that the adiabatic regime indeed extends from a learning rate of about 6e-8 (as was in Figure 2) to about 1.3e-7. Thus, freezing of output.dense.weight did help to somewhat extend the adiabatic tuning regime. However, further increase of the learning rate results in worse deterioration for the version with frozen output.dense.weight layer, as can be seen for XNLI starting from the rate 1.4e-7.
rate layer
1e-8 1.intermediate.dense.bias
3.intermediate.dense.bias
2e-8 5.intermediate.dense.weight
3.attention.output.LayerNorm.weight
3e-8 5.intermediate.dense.weight
1.attention.output.LayerNorm.weight
4e-8 5.intermediate.dense.weight
3.attention.output.LayerNorm.weight
5e-8 3.output.dense.weight
2.output.dense.weight
6e-8 3.output.dense.weight
2.output.dense.weight
7e-8 3.output.dense.weight
2.output.dense.weight
8e-8 1.output.dense.weight
3.output.dense.weight
9e-8 1.output.dense.weight
4.output.dense.weight
1e-7 1.output.dense.weight
5.output.dense.weight
Table 14: The ’most changed’ two layers at each learning rate. The ’change’ is defined as the maximal weight of the layer if it was changed by the tuning. The prefix ’encoder.layer’ is removed from the layer names here.
rate layer
1e-8 5.attention.output.dense.bias
11.output.dense.bias
2e-8 5.attention.output.dense.bias
11.output.dense.bias
3e-8 5.attention.output.dense.bias
11.output.dense.bias
4e-8 5.attention.output.dense.bias
11.output.dense.bias
5e-8 11.output.dense.weight
11.output.dense.bias
6e-8 11.output.dense.weight
11.output.dense.bias
7e-8 11.output.dense.weight
11.output.dense.bias
8e-8 11.output.dense.weight
11.output.dense.bias
9e-8 11.output.dense.weight
11.output.dense.bias
1e-7 11.output.dense.weight
11.output.dense.bias
Table 15: The ’most changed’ two layers at each learning rate. The ’change’ is defined as $(W_{t}-W_{o})/(W_{t}+W_{o})$, where $W_{t}$ is the maximal weight of the layer in the tuned query encoder, and $W_{o}$ is the maximal weight of the layer in the original (untuned) encoder. The prefix ’encoder.layer’ is removed from the layer names here.
M.4 Margin of triple loss
When using the triplet loss for contrastive learning, the margin is an important parameter that can significantly affect model training. In Figure 12 we show the dependency of the evaluation results on the margin during its tuning. We consider our default tuning parameters (Section 2.3), but change the margin. The results are not unexpected: a margin up to 0.15 is reasonable, and at higher margins the disturbance on cross-lingual, and, eventually, on English data evaluation becomes too strong.
Figure 12: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, learning rate 5e-8, batch size 14 using different triplet loss margins on (a) XNLI and (b) the English-only datasets (MSMARCO and ARXIV). Values that did not pass the two-tailed test are shown with open markers.
The corresponding data for $L12$ are given in Figure 13. It shows that a margin of 0.1 works best for $L12$. The results for margin 0.1 are distinctly better. Altogether, $L12$ appears to be more sensitive (compared to $E5$) to the tuning parameters if the goal is to preserve performance on multilingual XNLI data and on out-of-domain ARXIV data. Arguably, the margin value of approximately 0.1 is the best both for $L12$ and $E5$.
Figure 13: Evaluations of the $L12$ query encoder tuned with a frozen embedding block, learning rate 5e-8, batch size 14 using different triplet loss margins on (a) XNLI and (b) the English-only datasets (MSMARCO and ARXIV). Values that did not pass the two-tailed test are shown with open markers.
M.5 Stopping criterion
In Table 16 we show how the improvement depends on the stopping criterion. The stoppings after 5 or 10 non-improvement epochs give similar results. Stopping after 15 non-improvement epochs continues the trend of increased gain on English data, but with a deterioration on a few language pairs.
to stop cdcdcdc+/- d+/- c+/- d+/-
5 6.50 7.60 4.18 4.40 2.26 6.62 222/0 217/0 208/1 174/5
10 7.38 8.95 4.87 4.81 3.26 7.73 222/0 218/0 201/2 163/17
15 8.93 10.98 5.81 6.10 2.51 7.73 222/0 217/3 191/17 140/51
M.6 Execution time
There is no essential difference between the execution times for $E5$ and $L12$. The tuning time depends on how soon stopping happened. At the settings of interest (Section 2.33.13.2), the tuning on an A100 GPU takes about one hour. For example, tuning 10 times at the default settings (Section 2.3, Appendix E) for rates between 1e-8 and 1e-7 takes 9 hours. At higher rates, stopping occurs earlier; tuning 10 times for rates between 1.1e-7 to 2e-7 takes less than 5 hours. Table 1 (with freezing different parts of the encoder) was obtained in 6 hours.
Evaluation of an encoder on all datasets we used (MSMARCO, ARXIV-first, ARXIV-title and XNLI) takes about 1.2-1.3 hours.
M.7 Effects of learning rate scheduler and weight decay
msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
B Sch D cdcdcdc+/- d+/- c+/- d+/-
100 $Q$ 1.83 2.80 0.99 1.23 -0.26 0.96 200/1 172/3 69/71 43/110
$E_{0.98}$ $10^{-6}$0.41 -1.07 -0.29 -0.83 -0.26 -0.96 0/63 0/39 0/20 1/8
64 - - 1.58 2.28 0.78 1.07 0.78 1.20 178/1 156/3 72/50 48/87
$Q$ 0.05 0.19 -0.34 -0.59 -0.26 0.00 0/0 0/0 0/0 0/1
$E_{0.98}$ $10^{-6}$ 0.13 0.17 -0.23 -0.32 -0.26 0.00 0/0 0/0 0/0 0/1
$E_{0.95}$ $10^{-6}$0.75 -1.47 -0.62 -0.78 -0.52 -0.48 0/67 0/48 0/66 1/42
$E_{0.95}$ $10^{-5}$0.75 -1.47 -0.62 -0.78 -0.52 -0.48 0/67 0/48 0/66 1/42
- $10^{-4}$ 1.58 2.28 0.78 1.07 0.78 1.20 178/1 156/3 72/50 48/87
$L$ $10^{-4}$ 0.30 0.55 -0.31 -0.11 -0.78 -0.48 0/0 0/0 0/4 0/9
32 $Q$ $10^{-4}$ 0.12 -0.17 -0.29 -0.43 0.26 0.00 0/0 0/0 0/0 0/0
$L$ $10^{-4}$0.05 -0.21 -0.18 -0.19 0.78 0.00 0/0 0/0 0/0 0/0
$E_{0.98}$ $10^{-4}$ 0.27 0.46 -0.21 -0.16 0.52 0.24 26/0 39/0 0/0 0/0
16 $L$ 0.00 -0.15 -0.10 -0.56 0.78 0.24 0/0 0/0 4/0 10/0
$E_{0.95}$0.70 -1.21 -0.65 -0.78 0.26 -0.96 0/53 0/32 7/6 20/0
$E_{0.95}$ $10^{-4}$0.70 -1.21 -0.65 -0.78 0.26 -0.96 0/53 0/32 7/6 20/0
8 $Q$ $10^{-6}$0.81 -1.39 -0.57 -1.15 -1.30 -2.39 0/167 0/123 0/121 2/81
$E_{0.95}$ $10^{-6}$2.98 -4.31 -2.34 -2.73 -2.60 -6.22 0/220 0/208 2/187 15/128
- $10^{-5}$0.81 -1.43 -0.78 -1.18 -1.30 -2.87 0/165 0/126 0/115 3/68
$Q$ $10^{-4}$0.81 -1.39 -0.57 -1.15 -1.30 -2.39 0/167 0/123 0/121 2/81
$E_{0.95}$ $10^{-4}$2.98 -4.31 -2.34 -2.73 -2.60 -6.22 0/220 0/208 2/187 15/128
Table 17: Percentage improvement over the fine-tuned E5 model with a frozen embedding block and tuned using a batch size of $14$, learning rate $5e$-$8$ and a margin of $0.1$. The blue colors indicate the top improvements whereas the red colors indicate the worse degradation. Three parameters are randomly varied: the batch size (denoted as “B”), the learning rate scheduler (denoted as “Sch”) and the weight decay (denoted as “D”). The learning rate schedulers are defined in Table 19 with an initial learning rate of $5e$-$8$. cand drefer to measuring the similarity of the text pairs using either the cosine similarity or the euclidean distance, respectively. For XNLI, (+)(+)( + ) indicates the number of language pairs that were improved while ()(-)( - ) indicates those that have worsened out of a total of 225 language pairs. Note that only the statistically significant (determined by a Z-test) language pairs are retained and hence not all the improved/worsened counts sum to 225. Additionally, (ent-neutr) refers to entailment-entailment similarities compared with entailment-neutral similarities whereas (ent-contr) refers to comparisons against entailment-contradiction similarities.
@@ B Sch D cdcdcdc+/- d+/- c+/- d+/-
Using the fine-tuned $E5$ model with the frozen embedding block, tuned using a batch size of 14, and a margin of 0.1, we randomly vary the batch size, learning rate scheduler and weight decay in order to assess their impact on the model’s final performance. In Table 17 we present the change in performance across these different configurations for a learning rate of $5\times 10^{-8}$, which is our ‘default’ learning rate (Section 2.3). In Table 18 we do the same for a learning rate of $10^{-7}$. Table 19 lists the different schedulers we considered. Values in blue indicate the top improvements whereas values in red indicate the worse degradation.
Across these parameters, on average, the batch size appears to have the most significant impact, generally leading to poorer performance as the batch size is decreased. Within each batch size group, we see that using an exponential learning rate scheduler ($E_{0.95}$ or $E_{0.98}$) is generally worse than using any of the other schedulers or no scheduler at all. A specific exception exists when using a batch size of 100 where the exponential scheduler outperforms the quadratic one when the learning rate is set to $10^{-7}$. Across all the configurations considered, the most impact seems to be the one shown in the first row of Table 17, where we see good improvement over MSMARCO and ARXIV-first while simultaneously showing improvement over XNLI ent-neutr.
M.8 Varying the optimizer and learning rate
Table 20 shows the effects of choosing a different optimizer with a small and large learning rate. In addition to Adamax, we tried Adadelta and Stochastic Gradient Descent (SGD), both of which did not change the model weights in a significant enough way to affect the overall performance and hence, are not presented. For higher learning rates, SGD without momentum did elicit a change as shown in Fig. 14, but the trend in performance is similar to what is presented in Fig. 2 with higher resolution near the transition point between improved and degraded multilingual performance. At around $9\times 10^{-7}$, we see a sharp increase in the number of degraded language pairs while the model maintains constant improvement on MSMARCO. With a high enough learning rate, it seems that the gradients are able to overcome a barrier in the loss landscape that confined the weights to a region in which multilingual characteristics were preserved.
Figure 14: Evaluations on (a) XNLI and (b) the English-only datasets (MSMARCO and ARXIV) of the E5 query encoder tuned with a frozen embedding block, batch size 14, margin 0.1 using different learning rates. Here we tune using SGD without momentum. Values that did not pass the two-tailed test are shown with open markers.
From the table, the default version of Adamax (Adamax with momentum) has a nearly negligible effect on the model when used with a small learning rate, suggesting that this particular configuration for the optimizer forces the model weights to change very slowly. When momentum is switched off, the model weights change enough to improve the overall performance in both English and other languages. Continuing down to the bottom row, if we turn up the learning rate to a higher value, the model weights begin to change more significantly which brings about less improvement in the model’s multilingual capacity (still an improvement nonetheless), but maintains the same improvement on English. Overall, going from the first row to the last row (for Adamax), we transition from a point in model weight space where performance on all languages can be enhanced or preserved to a point which is better suited for the English-only task defined in tuning.
msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
O M LR cdcdcdc+/- d+/- c+/- d+/-
AdamW Yes 2e-8 6.50 7.62 4.63 4.60 2.76 5.96 222/0 218/0 208/1 171/5
AdamW Yes 1e-7 9.26 11.38 6.06 6.45 3.51 9.49 222/0 214/3 188/18 132/63
Adamax Yes 2e-8 0.81 1.14 0.62 0.63 0.75 -0.22 0/0 0/0 1/0 1/0
No 2e-8 6.40 7.50 4.25 4.63 3.01 6.40 224/0 220/0 214/1 178/4
Yes 1e-7 6.67 7.65 4.67 4.68 3.51 6.40 222/0 217/0 205/2 171/7
No 1e-7 6.46 7.60 4.67 5.06 3.51 6.40 222/0 217/0 198/2 157/12
Table 20: Percentage improvement over the untuned $E5$ model. O, M and LR represent the choice of optimizer, whether or not momentum was used and the learning rate, respectively. All the models here are tuned with a batch size of $14$, margin $0.1$, and a frozen embedding block. Adamax with no momentum corresponds to choosing $\beta_{1}=\beta_{2}=0$ for the optimizer parameters.